Grade upto college level Mechanics

Two mercury drops (each of radius ‘r’ ) merge to form bigger drop. The surface energy of the bigger drop, if T is the surface tension, is:
(a) 4πr2T
(b) 2πr2T
(c) 28/3πr2T
(d) 25/3πr2T

Amit Saxena
12 Years agoGrade upto college level

2 Answers

Navjyot Kalra
12 Years ago
(c) Sum of volumes of 2 smaller drops = Volume of the bigger drop
2.4/3πr3 = 4/3πR3 \RightarrowR = 21/3r
Surface energy = T.4πR2
= T4π2 r­2 = T.28/3 πr2
Kushagra Madhukar
5 Years ago
Dear student,
Please find your answer below.
 
Sum of volumes of 2 smaller drops = Volume of the bigger drop
2.4/3πr3 = 4/3πR3 \RightarrowR = 21/3r
Surface energy = T.4πR2
= T4π2 r­2 = 28/3 πr2T
Hence option c) is correct.
 
Regards,
Kushagra